OpenStudy (anonymous):

square root of 49 over (3+4i) - (2-5I)=

OpenStudy (anonymous):

What answer did you get?

OpenStudy (helder_edwin):

is it like this? \[\large \frac{49}{3+4i}-(2-51) \]

OpenStudy (anonymous):

sq root of 49 is over (3+4i) - (2-5i) please help!

OpenStudy (anonymous):

Nope. Like this: 49/(3+4i) - (2-51)

OpenStudy (anonymous):

Make sure your (2-51) is in the denominator as well with the (3+4i).

OpenStudy (helder_edwin):

so it is \[\large \frac{49}{(3+4i)-(2-5i)} \]

OpenStudy (anonymous):

Yes! Now, try to solve it and tell me your answer.

OpenStudy (helder_edwin):

just add or subtract like terms in the denominator

OpenStudy (anonymous):

Do you have a scientific calc?

OpenStudy (helder_edwin):

then use this \[\large \frac{z}{w}=\frac{z\overline{w}}{w\overline{w}} \] where \(\overline{w}=a-bi\) if \(w=a+bi\)

OpenStudy (anonymous):

I'm not quite sure, however, I'll just give you the answer. Answer: 49/82 - 441i/82